Shutdown pquest SNF 2009-05-07 16:46:29: RF system still not working

eenriquez at snf.stanford.edu eenriquez at snf.stanford.edu
Mon Jun 1 12:36:39 PDT 2009


RF now working.  Ran O2 plasma.  He flow at 10T is 7.2 sccm.  Cycled wafer in out of the system several times.  




More information about the pquest-pcs mailing list